OpenStudy (anonymous):

mnso4 hot to balance in oxidation no.

OpenStudy (chmvijay):

Mn+2 and SO4-2

OpenStudy (chmvijay):

[SO4]-2 mean O has -2 charge since 4O are therefore total charge on O is =-2*4=-8 since total charge outside the square bracket is -2 hence charge on sulfur should be +6

OpenStudy (chmvijay):

+6 -8=-2

OpenStudy (chmvijay):

Mn =+2 S=+6 O=-2